OpenStudy (anonymous):

A series of positive consective integers add upto 2000. Find the least value of m where m is the first term

OpenStudy (kainui):

Well if m is the first term, and these are consecutive numbers, then doesn't that mean the next term is (m+1)?

OpenStudy (anonymous):

yes

OpenStudy (kainui):

Wait, I think I misinterpreted this as being a really easy question when it's actually a little more involved. Hmm.

ganeshie8 (ganeshie8):

that should work right.. m + (m+1) + ... = 2000

OpenStudy (kainui):

Actually that's completely right. Just need to find a general summation formula.

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

say there are "n" terms in the series

ganeshie8 (ganeshie8):

m + (m+1) + ... + (m + n-1) = 2000

OpenStudy (anonymous):

well okay

ganeshie8 (ganeshie8):

nm + n(n-1)/2 = 2000

ganeshie8 (ganeshie8):

hmm guess we need to mess with this equation a bit..

OpenStudy (kainui):

@ganeshie8 Technically I think your equation should be: \[m(n+1)+\frac{n(n+1)}{2}=2000\]

ganeshie8 (ganeshie8):

yes if there are n+1 terms...

OpenStudy (kainui):

Wait I wrote it wrong and it happened to agree with what you wrote. What I mean is that it should agree with both.\[m(n+1)+\frac{n(n-1)}{2}=2000\]

OpenStudy (kainui):

Wait maybe I'm confusing myself, I thought one of your summations in n was off by 1 but now I'm not even sure. I think I need to go to bed now. Good luck, this is a very interesting problem. @No.name

ganeshie8 (ganeshie8):

n terms : (m+0) + (m+1) + ... + (m + n-1) = 2000 (m+m+... n times) + (0+1+ ... n-1)= 2000 (nm) + (n(n-1)/2) = 2000 but we still need to interpret this equation in integers and find the minimum value for m

OpenStudy (ikram002p):

there must be a trick about pascal number right ?

ganeshie8 (ganeshie8):

Is it correct to say that small \(m\) means large \(n\) ?

ganeshie8 (ganeshie8):

smallest first term => largest number of terms in the series

OpenStudy (ikram002p):

like 0+1+2+3+...+m+(m+1)+(m+2)+...+n = 2000 + m(m+1)/2

hartnn (hartnn):

i agree with that. smallest 1st term = largest no. of terms. also i am getting m=47 (n=32) by trial and error

ganeshie8 (ganeshie8):

m=47 is correct hartnn, i still don't have a solution though http://www.wolframalpha.com/input/?i=solve+n*m+%2B+n%28n-1%29%2F2+%3D+2000%2Cn%3E0%2Cm%3E0+over+integers

hartnn (hartnn):

(2m-1)^2 +16000 shold be a perfect square

ganeshie8 (ganeshie8):

Oh yes the quadratic gives that condition - but thats a necessary condition - i feel the solution is going to be complex than it appeared initially !

OpenStudy (ikram002p):

mmm isnt there a a condition to how sumation from 0 should be like or some form ?

ganeshie8 (ganeshie8):

1+2+3+... +n is always a triangular number, so ?

OpenStudy (ikram002p):

so , whats the nearest triangular number to 2000 ?

ganeshie8 (ganeshie8):

62(62+1)/2 = 1953

ganeshie8 (ganeshie8):

63(63+1)/2 = 2016

OpenStudy (ikram002p):

mmm whats the nearest triangular number that devide 3

OpenStudy (ikram002p):

mm forget it ...its trial method

ganeshie8 (ganeshie8):

Alright, lets work it using basic number theory

ganeshie8 (ganeshie8):

n*m + n(n-1)/2 = 2000 2nm + n(n-1) = 4000 n(2m + n - 1) = 4000

ganeshie8 (ganeshie8):

=> n | 4000 2m+n-1 | 4000

OpenStudy (ikram002p):

i was like trying to use this [0+1+2+3+...+(m-1) ]+[m+(m+1)+(m+2)+...+n] = m(m-1)/2+2000 and find triangular number (s) which is above 2000 such that s-2000 is also triangle number

OpenStudy (ikram002p):

oh , diophantine equation :O

ganeshie8 (ganeshie8):

yep

OpenStudy (ikram002p):

but to this point we have or n | 4000 2m+n-1 | 4000 it's not necessary that both divides 4000

ganeshie8 (ganeshie8):

ab = c => a|c b|c

OpenStudy (ikram002p):

hehe yeah xD sorry im just hungry hehehe ok continue plz

OpenStudy (ikram002p):

ok gtg nw , ill think of it ltr

OpenStudy (anonymous):

I don't have any such method but getting m= 47

hartnn (hartnn):

how ?

OpenStudy (anonymous):

I actually wanted a algebraic sollution mine is a more lengthier method , but i doon't know diaphante... or anything similar

OpenStudy (anonymous):

Just apply the Sum of terms in A.P formula , then you will get a quadratic when solve that then figure out some combinations

hartnn (hartnn):

oh, so trial and error...

OpenStudy (anonymous):

yeah you can say it trial and error , but is there any method

hartnn (hartnn):

not in a high school level math, if there is, even i haven't studied it

OpenStudy (ikram002p):

so @ganeshie8 lets continue from here n(2m + n - 1) = 4000

ganeshie8 (ganeshie8):

\[\large n(2m + n - 1) = 2^5\times 5^3\]

ganeshie8 (ganeshie8):

total # of factors = (5+1)(3+1) = 24

OpenStudy (ikram002p):

so n can take these values \(n=5\) \(n=5^2\) \(n=5^3\) \(n=2×5\) \(n=2×5^2\) \(n=2×5^3\) \(n=2^2×5\) \(n=2^2×5^2\) \(n=2^2×5^3\) \(n=2^3×5\) \(n=2^3×5^2\) \(n=2^3×5^3\) \(n=2^4×5\) \(n=2^4×5^2\) \(n=2^4×5^3\) \(n=2^5×5\) \(n=2^5×5^2\) \(n=2^5×5^3\) hehehe xD

OpenStudy (ikram002p):

yeah lol ! got it , continue

ganeshie8 (ganeshie8):

you forgot one value, n can take \(1\) too

OpenStudy (dan815):

u can do minimizing function right

OpenStudy (dan815):

callcooolus

OpenStudy (ikram002p):

will no since if n=1 then m=2000

ganeshie8 (ganeshie8):

callcoolus is not so cool when it comes to solving in integers danny

OpenStudy (dan815):

ohh right :)

OpenStudy (ikram002p):

im kind of stupid with calcules xD

ganeshie8 (ganeshie8):

ikram, did u get how we narrowed down to 2 cases ?

OpenStudy (ikram002p):

not realy

ganeshie8 (ganeshie8):

\[\large n(2m + n - 1) = 2^5\times 5^3\] n = 2^5 or n = 5^3 we can ignore all other cases cuz n and (2m+n-1) can not be both even / both odd at the same time.

OpenStudy (ikram002p):

why ?

ganeshie8 (ganeshie8):

if n is even, then (2m+n-1) has to be odd, right ?

OpenStudy (ikram002p):

right ?

ganeshie8 (ganeshie8):

if n is odd, then (2m+n-1) has to be even, right ?

OpenStudy (ikram002p):

ok ?

ganeshie8 (ganeshie8):

that means, n and (2m+n-1) cannot share the factor "2", right ?

OpenStudy (ikram002p):

aha ok

OpenStudy (kainui):

Maybe since there are two unknowns, we just need to create one more equation. Perhaps since there's a kind of symmetry we can come up with an equation for the average of the numbers.

OpenStudy (kainui):

Why aren't there nonuniform operators in math? It's always every term is exactly 1 term greater than the next or division is always into equal parts instead of unequal parts. If we could just divide this into just slightly less equal parts we could get our answer real easily lol.

OpenStudy (ikram002p):

but then mmm more cases could be , n =5,5^2,5^3 why u only chosed 5^3 also n=2,2^2....

OpenStudy (kainui):

Well the average distance between points always has the average equal to the middle distance if there's an odd number since the one on the left is one less than the center and the one on the right has one more so they cancel out. Does this make sense or am I not really doing anything useful here.